# Instructions for *ACL Proceedings

### v1

@@ Abstract
Instructions for *ACL Proceedings
Anonymous ACL submission
@@ 1 Introduction
In modern enterprises, business rules codify essential policies and procedures that govern automated decision-making in domains such as finance, healthcare, and e-commerce nelson2014business; yu2013modeling; kung2007fruits. Traditional rule engines and decision-management systems require domain experts to manually specify these rules and their execution logic, a process that is well-known to be labor-intensive and error-prone sivasankari2020comparative. Consequently, organizations are increasingly seeking to automate routine processes and decisions to improve efficiency and productivity. Nevertheless, automating complex, interdependent business logic remains a challenging problem.
Despite advances in natural language processing (NLP) for procedural text, existing datasets and models friedrich2011process; epure2015automatic; quishpi2020extracting; qian2020approach; lopez2021declarative; DBLP:conf/acl/LiangLDJLW023; DBLP:conf/acl/RenZ0ZL23; du2024paged focus primarily on action sequences, where systems extract sequences of tasks or operations (e.g., "chop onions" → "heat oil" → "fry onions" in a recipe). These approaches-typified by datasets like PAGED dataset du2024paged or works on procedural graph extraction-model processes as chains of actions with temporal or logical transitions. However, business processes often hinge not on explicit sequences of actions, but on interdependent rules, which govern decision logic under varying conditions. For instance, in a banking scenario, the rule ‘If the currency type is USD, then select remittance type’ depends on a condition being met, rather than an explicit preceding action. Further decisions, such as verifying identity or collecting additional input, may occur in parallel or be conditionally triggered based on earlier rule outcomes. This distinction between action-based and rule-based modeling is crucial: action flows capture what is done in sequence, while rule flows express under what conditions actions occur. Yet, existing procedural modeling works largely overlook this rule-centric structure, leaving a gap in automated understanding of real-world business logic.
@@ 2.1 Procedural Graph Extraction
Prior work on procedural graph construction has mainly focused on modeling action sequences in process-oriented texts. For instance, maeta2015framework applied named entity recognition techniques to extract procedural concepts from texts (e.g., recipes), and constructed process graphs by identifying action-object, tool-action, and state-transition dependencies. guo2018automatic introduced a method tailored to Chinese emergency response plans, leveraging structured text formats and syntactic rules to extract response conditions, roles, and task dependencies, thereby generating BPMN models. Similarly, ren2023constructing proposed a framework to extract multiple types of dependencies, including sequential, inclusion, and constraint, from unstructured procedural documents to construct enriched procedural graphs. pal2021constructing combined pre-trained language models with graph neural networks (LM-GNN) to identify action-information flows from unstructured texts, while candido2024annotated employed BiLSTM+CRF to extract entities and their interrelations from business process descriptions for model generation.
@@ 2.2 LLMs for Business Process Modeling
Recent advances in LLMs have demonstrated significant potential across a range of NLP tasks, including business process modeling. bellan2022extracting were among the first to explore LLMs usage for extracting business process elements, though their approach was primarily limited to identifying sequential dependencies and lacked mechanisms to capture complex structures such as parallelism and conditional branching. du2024paged introduced a large-scale dataset for process reasoning and fine-tuned LLMs within commercial domains (e.g., restaurant services), yet their findings remain narrowly scoped, with limited generalizability to domains such as healthcare and manufacturing.
@@ 3 BPRF Dataset
[…]
@@ 3.2 Dataset Construction
Our dataset is derived from both real business environments, including banks, e-commerce platforms, flight booking system, and other relevant domains. Additionally, we incorporate synthetic data generated via ChatGPT DBLP:conf/nips/Ouyang0JAWMZASR22. The dataset construction process is shown in Fig 1, involving the following steps:
  1. [leftmargin=*, itemindent=0pt]
@@ 3.3 Dataset Analysis
Human Evaluation of Business Text Quality: Following prior studies miller1979humanistic; DBLP:conf/acl/DuLLL24, we engaged three annotators to assess generated business texts across five key dimensions: Readability, Accuracy, Clarity, Simplicity and Usability. Annotators rated each criterion on a scale from 1 to 5. The evaluation results across these dimensions are depicted in Fig 2. To assess reliability of annotations, we computed Intraclass Correlation Coefficient (ICC) shrout1979intraclass among annotators, with a higher ICC indicating stronger inter-annotator agreement and more reliable annotations. An ICC value of 0.75 or higher is generally considered indicative of strong agreement koo2016guideline. Finally, we calculated the average ICC across all criteria, obtaining a value of 0.904, which indicates strong agreement among annotators.
Inter-Annotator Agreement for Business Rule Annotation: Standard inter-annotator agreement (IAA) artstein2017inter measures primarily focus on classification tasks or qualitative numerical annotations, making it challenging to evaluate datasets annotated with structured business rules. To address this, we transformed our dataset into a Named Entity Recognition (NER) format, allowing for standard agreement measures. An example annotation is provided in Table 3. Three NLP experts in the phase of Verification annotated the dataset using the BIO tagging scheme. We then computed Fleiss’ Kappa artstein2017inter to measure inter-annotator agreement, obtaining a score of 0.911, which indicates almost perfect agreement (values range from 0 to 1, with scores above 0.6 considered high).
@@ 4 Methodology
In this section, we propose ExIde, a two-stage framework for automated business process extraction using LLMs, as illustrated in Fig 3. The framework includes: (1) Extraction of Business Rules and (2) Identification of Dependency Relationships. In the first stage, we design five prompt variants to guide the LLMs in extracting business rules from domain-specific texts. These prompts adopt a chain-of-thought format wei2022chain, structured according to a shared template shown in Fig 8, where Definition 1, 2, 4 refer to the definitions introduced in Section 3.1. The specific details of five prompts are as follows:
  • [leftmargin=*, itemindent=0pt]
@@ 5 Experiments
In this section, we present a comprehensive evaluation of LLMs for task of automatic business process extraction from our curated dataset. Specifically, we assessed 12 representative LLMs, including 5 closed-source models: GPT 4o hurst2024gpt, GPT o1 o1, ERNIE3.5 ernie3.5, QWEN2.5 qwen, and ChatGLM glm2024chatglm, where these models are based on web versions. The open-source group comprises QWEN 2.5-14b qwen2.5, Llama 3.1-8b llama, DeepSeek v3 deepseekai2024deepseekv3technicalreport, DeepSeek r1 guo2025deepseek, ERNIE-3.5-128K, GLM-4-9b-chat and QWQ-32b qwq32b. Notably, GPT o1, DeepSeek r1, and QWQ-32b are explicitly designed with advanced reasoning capabilities.

### v2

# Business as Rulesual: A Benchmark and Framework for Business Rule Flow Modeling with LLMs

@@ Abstract
Business as Rulesual: A Benchmark and Framework for Business Rule Flow Modeling with LLMs
Chen Yang1, Ruping Xu1, Ruizhe Li2, Bin Cao1thanks: Corresponding author., Jing Fan1 Zhejiang University of Technology, China1 University of Aberdeen, UK 2 {yangchen,rupingxu,bincao,fanjing}@zjut.edu.cn1 ruizhe.li@abdn.ac.uk2
@@ 1 Introduction
In modern enterprises, business rules codify essential policies and procedures that govern automated decision-making in domains such as finance, healthcare, and e-commerce Nelson and Sen (2014); Yu et al. (2013); Küng and Hagen (2007). Traditional rule engines and decision-management systems require domain experts to manually specify these rules and their execution logic, a process that is well-known to be labor-intensive and error-prone Sivasankari et al. (2020). Consequently, organizations are increasingly seeking to automate routine processes and decisions to improve efficiency and productivity. Nevertheless, automating complex, interdependent business logic remains a challenging problem.
Despite advances in natural language processing (NLP) for procedural text, existing datasets and models Friedrich et al. (2011); Epure et al. (2015); Quishpi et al. (2020); Qian et al. (2020); López et al. (2021); Liang et al. (2023); Ren et al. (2023a); Du et al. (2024a) focus primarily on action sequences, where systems extract sequences of tasks or operations (e.g., "chop onions" → "heat oil" → "fry onions" in a recipe). These approaches-typified by datasets like PAGED dataset Du et al. (2024a) or works on procedural graph extraction-model processes as chains of actions with temporal or logical transitions. However, business processes often hinge not on explicit sequences of actions, but on interdependent rules, which govern decision logic under varying conditions. For instance, in a banking scenario, the rule ‘If the currency type is USD, then select remittance type’ depends on a condition being met, rather than an explicit preceding action. Further decisions, such as verifying identity or collecting additional input, may occur in parallel or be conditionally triggered based on earlier rule outcomes. This distinction between action-based and rule-based modeling is crucial: action flows capture what is done in sequence, while rule flows express under what conditions actions occur. Yet, existing procedural modeling works largely overlook this rule-centric structure, leaving a gap in automated understanding of real-world business logic.
@@ 2.1 Procedural Graph Extraction
Prior work on procedural graph construction has mainly focused on modeling action sequences in process-oriented texts. For instance, Maeta et al. (2015) applied named entity recognition techniques to extract procedural concepts from texts (e.g., recipes), and constructed process graphs by identifying action-object, tool-action, and state-transition dependencies. Guo et al. (2018) introduced a method tailored to Chinese emergency response plans, leveraging structured text formats and syntactic rules to extract response conditions, roles, and task dependencies, thereby generating BPMN models. Similarly, Ren et al. (2023b) proposed a framework to extract multiple types of dependencies, including sequential, inclusion, and constraint, from unstructured procedural documents to construct enriched procedural graphs. Pal et al. (2021) combined pre-trained language models with graph neural networks (LM-GNN) to identify action-information flows from unstructured texts, while Candido et al. (2024) employed BiLSTM+CRF to extract entities and their interrelations from business process descriptions for model generation.
@@ 2.2 LLMs for Business Process Modeling
Recent advances in LLMs have demonstrated significant potential across a range of NLP tasks, including business process modeling. Bellan et al. (2022) were among the first to explore LLMs usage for extracting business process elements, though their approach was primarily limited to identifying sequential dependencies and lacked mechanisms to capture complex structures such as parallelism and conditional branching. Du et al. (2024a) introduced a large-scale dataset for process reasoning and fine-tuned LLMs within commercial domains (e.g., restaurant services), yet their findings remain narrowly scoped, with limited generalizability to domains such as healthcare and manufacturing.
@@ 3 BPRF Dataset
Our bank supports up to 39 currency types of popular countries or regions around the world, including RMB, USD, JPY, GBP, and HKD. After selecting the appropriate currency, the customer needs to choose the corresponding cash or remittance type based on the type of business to be conducted thereafter. The cash or remittance type includes cash and remittance. Upon completing the cash/remittance type selection, the customer needs to provide the purchase amount
Slot type Logical judgement Reference value Action
@@ 3.2 Dataset Construction
Our dataset is derived from both real business environments, including banks, e-commerce platforms, flight booking system, and other relevant domains. Additionally, we incorporate synthetic data generated via ChatGPT Ouyang et al. (2022). The dataset construction process is shown in Fig 1, involving the following steps:
@@ 3.3 Dataset Analysis
Human Evaluation of Business Text Quality: Following prior studies Miller (1979); Du et al. (2024b), we engaged three annotators to assess generated business texts across five key dimensions: Readability, Accuracy, Clarity, Simplicity and Usability. Annotators rated each criterion on a scale from 1 to 5. The evaluation results across these dimensions are depicted in Fig 2. To assess reliability of annotations, we computed Intraclass Correlation Coefficient (ICC) Shrout and Fleiss (1979) among annotators, with a higher ICC indicating stronger inter-annotator agreement and more reliable annotations. An ICC value of 0.75 or higher is generally considered indicative of strong agreement Koo and Li (2016). Finally, we calculated the average ICC across all criteria, obtaining a value of 0.904, which indicates strong agreement among annotators.
Inter-Annotator Agreement for Business Rule Annotation: Standard inter-annotator agreement (IAA) Artstein (2017) measures primarily focus on classification tasks or qualitative numerical annotations, making it challenging to evaluate datasets annotated with structured business rules. To address this, we transformed our dataset into a Named Entity Recognition (NER) format, allowing for standard agreement measures. An example annotation is provided in Table 3. Three NLP experts in the phase of Verification annotated the dataset using the BIO tagging scheme. We then computed Fleiss’ Kappa Artstein (2017) to measure inter-annotator agreement, obtaining a score of 0.911, which indicates almost perfect agreement (values range from 0 to 1, with scores above 0.6 considered high).
@@ 4 Methodology
In this section, we propose ExIde, a two-stage framework for automated business process extraction using LLMs, as illustrated in Fig 3. The framework includes: (1) Extraction of Business Rules and (2) Identification of Dependency Relationships. In the first stage, we design five prompt variants to guide the LLMs in extracting business rules from domain-specific texts. These prompts adopt a chain-of-thought format Wei et al. (2022), structured according to a shared template shown in Fig 8, where Definition 1, 2, 4 refer to the definitions introduced in Section 3.1. The specific details of five prompts are as follows:
@@ 5 Experiments
In this section, we present a comprehensive evaluation of LLMs for task of automatic business process extraction from our curated dataset. Specifically, we assessed 12 representative LLMs, including 5 closed-source models: GPT 4o Hurst et al. (2024), GPT o1 OpenAI (2024), ERNIE3.5 Team (2023), QWEN2.5 Team (2024b), and ChatGLM GLM et al. (2024), where these models are based on web versions. The open-source group comprises QWEN 2.5-14b Team (2024a), Llama 3.1-8b AI (2024), DeepSeek v3 DeepSeek-AI and Aixin Liu (2024), DeepSeek r1 Guo et al. (2025), ERNIE-3.5-128K, GLM-4-9b-chat and QWQ-32b Team (2025). Notably, GPT o1, DeepSeek r1, and QWQ-32b are explicitly designed with advanced reasoning capabilities.
@@
6 Conclusion
In this paper, we introduced a novel framework, ExIde, for extracting business rules and their interdependencies from unstructured business process documents using LLMs. Moreover, we constructed a comprehensive Chinese dataset, BPRF, containing 326 labeled business rules across diverse domains. Our experiments demonstrated the effectiveness of ExIde in accurately extracting business rules and identifying their logical dependencies using 12 state-of-the-art LLMs. Notably, the results highlight the importance of prompt design and reasoning-oriented model architectures in enhancing extraction accuracy. By bridging the gap between rule-centric business logic and LLMs, our work paves the way for more intelligent, interpretable, and automated business process modeling systems. Future work includes expanding the dataset, refining prompt strategies, and integrating rule execution into downstream decision support and automation systems.
@@ 6 Conclusion
Despite the promising results demonstrated by ExIde and the contributions of the BPRF dataset, several limitations remain. First, while our dataset spans multiple domains (e.g., banking, healthcare, e-commerce), its size—326 business rules across 50 documents—may still be insufficient to capture the full diversity and complexity of business process logic in the real world. This could affect the generalizability of our findings. Second, a portion of our dataset is synthetically generated using ChatGPT, which, while enhancing coverage and diversity, may introduce distributional biases not representative of naturally occurring business texts. Third, our framework currently relies on prompt-based querying of LLMs, which may suffer from sensitivity to prompt formulation and inconsistencies across different model architectures and decoding parameters. Lastly, the current approach focuses primarily on extracting and classifying pairwise dependencies; modeling more complex global structures or constraints remains an open challenge.
Ethical Statement
This work presents a novel benchmark (BPRF) and framework (ExIde) for extracting business rules and their logical dependencies from business process documents. The dataset includes both real-world business texts collected from publicly available sources and synthetically generated texts using ChatGPT. All real data were obtained from publicly accessible documents and do not contain any personally identifiable information (PII) or confidential content. Moreover, no human subjects were involved in the model evaluation phases, and no sensitive user data were used. The study adheres to responsible AI development principles, including transparency, reproducibility, and minimization of potential harms. Our annotated dataset will be released upon paper acceptance to support further research in this area.